What is the ratio of lactic acid (Ka = 1.37x10-4) to lactate in a solution with pH =4.36?
1
Expert's answer
2021-10-01T13:32:02-0400

"pH=pka +Log" [ (lactate ion) / (lactic acid ) ]

pH = 4.36 , "pka" = -log "1.37 \u00d7" 10-4


=3.86


4.36=3.86 + Log [ (lactate ion) / ( lactate acid ) ]

Log [ lactic ion / lactic acid ] =4.36 -3.86 = 0.50


[ lactic ion / lactic acid ] = 100.50 = 1.1220 / 1


Ratio of lactic acid to lactate Ion


= 1 / 1.1220


=1.1220 : 1